I am to recall that, when interest is compounded continuously, the balance in an account after t years is given by
A = Pe^rt,
where P is the initial investment and r is the interest rate.
A computer is infected with the Sasser virus. Assume that it infects 20 other computers within 5 minutes; and that these PCs and servers each infect 20 more machines within another five minutes, etc. How long until 100 million computers are infected?
Thank you very much!

Answer by stanbon(48535)

I am to recall that, when interest is compounded continuously, the balance in an account after t years is given by A = Pe^rt, where P is the initial investment and r is the interest rate.
--------------------------------
A computer is infected with the Sasser virus. Assume that it infects 20 other computers within 5 minutes; and that these PCs and servers each infect 20 more machines within another five minutes, etc. How long until 100 million computers are infected?
---------------------------------
100x10^6 = 20^(t/5)
Take the log to get:
(t/5)log20 = 8
t = 40/log(20)
t = 30.745.. minutes
=====================
Cheers,
Stan H.